# Balanced Resonate-and-Fire Neurons

### v1

@@ 1 Introduction
Artificial neural networks (ANNs) have become the main method for solving machine learning problems in recent years (Goodfellow et al., 2016; Wu & Feng, 2018; Abiodun et al., 2019). However, they require massive computation and energy, making them inefficient for large-scale real-world applications—particularly in the realm of edge computing—due to the deep non-linear continuous estimators to represent features of the data (Pfeiffer & Pfeil, 2018). Spiking neural networks (SNNs) circumvent this drawback by processing information through the precise timing of action potentials, or spikes. SNNs can potentially be more efficient compared to ANNs due to their event-driven property, for which computation is only required when spikes are propagated in the system (Paugam-Moisy & Bohté, 2012). Furthermore, SNNs are self-recurrent, as the neurons have a dynamic internal state that modulates their activity over time, and are more biologically realistic than conventional ANNs. Some examples of spiking neurons with ascending biological plausibility include the leaky integrate-and-fire (LIF) neuron, the adaptive leaky integrate-and-fire (ALIF) neuron (Bellec et al., 2018), the Izhikevich neuron (Izhikevich, 2003), and the Hodgkin-Huxley (HH) neuron (Hodgkin & Huxley, 1952). While the HH model is too computationally expensive for practical application, it exhibits rich and complex membrane dynamics that simpler models lack.
@@
One mode of dynamics overlooked in ML applications in particular is the resonating behavior seen in oscillatory neurons. Subthreshold oscillations of the membrane potential have been observed in various mammalian nervous systems, including the neurons in the frontal cortex (Llinas et al., 1991), in the thalamus (Pedroarena & Llinás, 1997), and in layer II of the medial entorhinal cortex (Alonso & Llinás, 1989), which is crucial in spatial information processing (Doeller et al., 2010). The resonate-and-fire (RF) neuron proposed by Izhikevich (2001) models such damped or sustained subthreshold oscillations of biological neurons: the RF neuron fires when the frequency of the incoming spikes matches that of the neuron’s damped oscillation. Incoming signals with higher frequency leads to more firing for a LIF neuron but less firing for an RF neuron with slow oscillation. As the RF neuron is similarly computationally efficient as the LIF neuron, it is potentially a spiking neuron model suitable for large-scale SNNs (Izhikevich, 2001).
@@
Previous works showed RF neurons implemented in the Intel Loihi 2 (Davies et al., 2018), a neuromorphic processor, can be used to compute the short-time Fourier transform (STFT) of signals (Frady et al., 2022) more computationally efficient than conventional STFT (Shrestha et al., 2023). Moreover, the RF neurons successfully converted raw data signals to spike trains, further input to SNNs with LIF neurons for detection and classification tasks (Hille et al., 2022; Lehmann et al., 2023). RF neurons have also been implemented in the framework of (feedforward) SNNs as harmonic oscillators for image classification tasks (AlKhamissi et al., 2021), as well as in optical flow estimation and audio classification tasks (Frady et al., 2022). Still, the performance of the RF models did not significantly exceed that of deep LIF SNNs (Frady et al., 2022) or LSTM cells (AlKhamissi et al., 2021), and these studies did not compare with state-of-art spiking neuron models, such as ALIF. Moreover, they lack of meaningful parameter studies and did not address stability issues of the vanilla RF model. Considering the strengths of the RF neuron, these results suggest that the full potential of the RF neuron has yet to be explored.
@@ 2 Resonate-and-Fire Neurons
Refer to captionptunstable($\omega{=}10$, $b{=}-0.3$)ptstable($\omega{=}10$, $b{=}-1$)
@@ 3 Balanced RF Models
The influence of implementing the refractory period and the smooth reset can be seen in Figure 2. The single RF neuron was simulated with $\omega=10,\,b=-1$, and the input signal was frequency-timed. The combination of refractory period and smooth reset significantly reduced the number of spikes, while the output spikes effectively reflected the period of the angular frequency of the neuron. Without both, the RF neuron was not able to release single spikes but tended to burst.
@@
To alleviate the divergence problem, we propose to restrict the parameter space to a subspace below the divergence boundary—an analytically derived relationship between $\delta$, $b_{c}$, and $\omega$ that ensures convergence. For an RF neuron to converge or show sustained oscillation, the magnitude of the membrane potential decreases or stays constant with time given an incoming signal:
@@
Furthermore, the explicit form of the subthreshold oscillatory behavior of the membrane potential after an incoming spike is derived in Appendix A.3 and given as:
@@
With a default $\delta=0.01$, the highest frequency that the neuron can resonate at is the frequency corresponding to an angular frequency of $100~{}rad/s$. We define this angular frequency as the upper boundary $\omega_{ub}$. The upper bound of $b_{c}$ that leads to sustained oscillation is implemented in this paper as $p(\omega)$, also considered the divergence boundary:
@@
combined with a trainable b-offset $b^{\prime}>0$ to ensure flexibility and convergence $b_{c}=p(\omega)-b^{\prime}$ which is constant throughout one sequence length. Figure 9 shows exemplary divergence boundaries for various $\delta$ values. By combining the derived dampening factor with the smooth reset, we present the final equation of $b(t)$ used in this paper:
@@
Implementing the refractory period, smooth reset, and the divergence boundary to efficient learning and sparse spiking for all four datasets.
@@
The frequency response plots shown in Figure 3 assess how well the RF neuron responds to specific frequencies. A peak on a frequency response plot indicates a high sensitivity of the neuron to the input signal frequency (details in Appendix A.4). The figure shows alignment of the response peak and $\omega$ of the RF neuron, demonstrating the resonance property to also be present in the discrete case. The slight offset of the response peak and $\omega$ of the RF neuron is an artifact of the numerical integration of the membrane dynamics’ differential equation and disappears with decreasing the temporal step size $\delta$.
@@
The RF neuron can also be considered a narrow band pass filter, for which only a specific range of frequencies are filtered depending on its angular frequency. Combined with the discrete time scale and b-offset, they determine how narrow and sensitive the band pass filter is. At a global scale, the RF neuron can also be considered a low pass filter, as it can only filter frequencies below the corresponding angular frequency $\omega_{ub}$ (Equation 19).
@@ 4 Network Implementation
We implement the BRF and BHRF neurons into the RSNN and applied them to simulations with several benchmark datasets. The essential formulations of the BRF and BHRF neuron are summarized in Algorithm 1 and Algorithm 2, respectively. Note that for the algorithmic formulation we change the notation to time-discrete tensor operations and superscript the time index, whereas the transition from $t$ to $t+1$ is considered as a time delay of $\delta$.
@@
We trained the networks with BPTT with the double-Gaussian function (Equation 38, Yin et al., 2021) as the surrogate gradient (Neftci et al., 2019). Further details of the networks are described in Appendix A.6.
Figure 4: Overview over the datasets. Examplary MNIST image and its corresponding sequential and permuted representations. Common pixel row outlined in red on MNIST and S-MNIST sample. ECG sample after level cross encoding. SHD sample after preprocessing.
@@ 5 Experiments
The MNIST dataset consists of grayscale 28$\times$28 pixel hand-written digit images for classification. The sequential-MNIST (S-MNIST), which converts the image to a sequence of 1 $\times$ 784, is a prominent benchmark dataset that enable comparison between sequential models with 54,000 images for training, 6,000 for validation, and 10,000 for testing. In the permuted S-MNIST variant (PS-MNIST), the pixel positions were first permuted randomly and then kept fixed.
Electrocardiogram (ECG) recordings of the heart are represented in voltage over time and consist of cyclic activity with six characteristic waveforms: P, PQ, QR, RS, ST, and TP. The QT database consisted of ECG recordings with per-time step labels classified by experts in the field (Laguna et al., 1997). There were 105 recordings, each recorded for 15 minutes with two electrodes. We left out 24 signals, as they were sudden death recordings without annotation. For the preprocessing shown in Figure 4, refer to Section A.7 in the appendix.
@@ 6 Results
Task Model Architecture No. params. (normal-$\downarrow$) Test Acc. (normal-$\uparrow$) SOPs (normal-$\downarrow$) SOPs/step (normal-$\downarrow$) SOP Ratio (normal-$\downarrow$)
@@
RF 1,256,10 68,874 98.0±$\pm 0.4~{}\%$ 29,034.8 37.03 0.41$\times$
BRF 1,256,10 68,874 99.0±$\pm 0.1~{}\%$ 15,462.6 19.72 0.22$\times$
BHRF 1,256,10 68,874 99.1±$\pm 0.1~{}\%$ 21,565.7 25.51 0.30$\times$
@@
RF 4,36,6 1,734 85.4±$\pm 0.5~{}\%$ 11,966.9 9.21 0.34$\times$
BRF 4,36,6 1,734 85.8±$\pm 0.7~{}\%$ 6,307.7 4.85 0.18$\times$
BHRF 4,36,6 1,734 87.0±$\pm 0.4~{}\%$ 6,233.8 4.80 0.18$\times$
SHD DCLS-D** 700,$(256)^{2}$,20 $\approx$200,000 95.1±$\pm 0.2~{}\%$
@@
2-8[2pt/1pt]
ALIF 700,$(128)^{2}$,20 142,120 90.4 % 24,690.0 98.76 1.00$\times$
RF 700,128,20 108,820 90.4±$\pm 0.7~{}\%$ 4,691.7 18.77 0.19$\times$
BRF 700,128,20 108,820 91.7±$\pm 0.8~{}\%$ 3,502.6 14.01 0.14$\times$
BHRF 700,128,20 108,820 92.7±$\pm 0.7~{}\%$ 4,139.5 16.56 0.17$\times$
The results, the network architecture, the number of parameters, and the number of produced spikes, between the best-performing BRF- and BHRF-RSNN and the baseline ALIF-RSNNs (see Section A.2) or other state-of-the-art (SoTa) models, are compared for each classification task in Table 1. The BHRF and BRF models outperformed the RSNN SoTa for three and two tasks respectively. Complementing accuracy, we also find that much enhanced sparseness in the BRF and BHRF networks while using substantially fewer parameters compared to the ALIF networks.
We also compare the theoretical energy efficiency of the RF and ALIF models as by computing their respective total spiking operations (SOPs) (Equation 40) and SOPs per sequence step. As both measures were considerably smaller for the BRF and BHRF models, this implies they require less computation to achieve better or comparable performance. This may be due to the resonating spiking behavior of the neuron, needing fewer spikes to represent its frequency, as well as the refractory period and smooth reset that hinders continuous spiking. A particularly large difference was seen in the SHD task, where the BRF model spiked only 14% of the ALIF model. Compared to the standard RF model, our balanced variants achieved significantly sparser activity while obtaining higher performance. Especially for the PS-MNIST, the standard RF model stayed at chance, while our balanced RF variations outperformed RSNN-SoTA.
Overall, we find that RSNNs comprised of either BRF and BHRF neurons exceeded performance for all classification tasks compared to the baseline ALIF-RSNN. Compared to standard RF neurons, the smooth reset, refractory period, and divergence boundary significantly improved the stability and learning of the RF parameters, possibly exploiting the model’s resonant properties. We study this in detail below.
@@
Task Model $\mathbf{95\,\%}$ (normal-$\downarrow$) $\mathbf{98\,\%}$ (normal-$\downarrow$) $\mathbf{100\,\%}$ (normal-$\downarrow$)
S-MNIST
ALIF
105 162 276
BRF
3 29 246
BHRF
5 12 119
PS-MNIST
ALIF
143 200 265
BRF
10 39 282
BHRF
6 19 123
ECG
ALIF
51 157 282
BRF
5 31 75
BHRF
15 38 109
SHD
ALIF
14 15 16
BRF
2 3 7
BHRF
3 5 8
@@
Figure 5: Top row: S-MNIST, PS-MNIST, ECG, and SHD learning curve between BRF and ALIF model. Each curve averaged per epoch (solid line) with standard deviation (shaded area) over five runs. The dot on the accuracy curves depict the point at which 95 % of the final accuracy was reached. Bottom row: S-MNIST, PS-MNIST, ECG, and SHD initial and optimized BRF parameter combinations: angular frequency $\omega$ and b-offset $b^{\prime}$ for all runs. Dashed line is the divergence boundary.
The learning curves of the BRF and ALIF models are presented in Figure 5. We reduced the ALIF model to the same size and number of trainable parameters as the BRF model. For the S-MNIST and PS-MNIST tasks, the ALIF networks only learned effectively when the truncated BPTT (TBPTT) with truncation step 50 was applied. Nonetheless, the fully back-propagated BRF networks converged significantly faster than the TBPTT ALIF networks, also evident from the quantitative convergence results (Table 2). The BRF networks showed especially stable convergence, as the five runs converged with only a small standard deviation. A similar pattern of fast and stable convergence was observed for the ECG dataset and SHD.
While numerically large, the learning rates of the BRF-RSNNs led to stable and fast convergence for most datasets. This is in contrast to standard SNNs where large learning rates lead to poor performance. Here, the double-Gaussian surrogate gradient function (Yin et al., 2021) we used may have contributed by ensuring gradient flow even for neurons far from firing. Combined with the high learning rate, this may have enabled non-spiking neurons to effectively shift their angular frequencies.
The ECG task was learned most effectively for the BRF and BHRF networks with a small batch size and high learning rate. A smaller batch size captures detailed variations within the dataset, notably the timing of the input spikes, which may have been important for per-step wave classification.
@@
For the S-MNIST task, plotting $\omega$-$b_{c}$ (Figure 5) shows clusters around $\omega$ values of 18 to 28 $rad/s$ and around $\omega$ of 40 $rad/s$ with a wide range of $b_{c}$. The dampening factor close to zero indicates near-sustained oscillation behavior, preserving resonance amplitude for an extended duration. The clustered variety in $b_{c}$ values suggests that short- and long-term memory are essential for effective learning of the S-MNIST task in B(H)RF networks. Additionally, the distribution of $\omega$ for S-MNIST task showed a bimodal distribution with the highest peak at 23.3 to 25.3 $rad/s$ and the second peak at 40.8 to 42.8 $rad/s$. The MNIST digits all contain some form of diagonal line with a width of 3 to 6 pixels, most prominently observed in digit 1. When converted into a sequence row-by-row, the lines are distributed among the sequence with a periodic signal of about 25 pixels. Considering the pixels as discrete time steps of $\delta$ = 0.01, the theoretical period T frequently observed in the signal is T = 0.25 $s$. Thus, the most frequent theoretical angular frequency $\omega^{\prime}$ underlying the signal is around $\omega^{\prime}=2\pi/0.25\,s=25.13~{}rad/s$ which is close to the most frequently learned angular frequencies of the BRF neuron. This simple calculation thus suggests that the BRF neurons were able to learn meaningful frequencies that underlie the S-MNIST dataset. For the PS-MNIST, we see more spread of $\omega$ and $b_{c}$ parameters, as the specific optimal frequency in S-MNIST is removed by the permutation.
Inspecting spike train ECG samples (Figure 4), we observe distinguishable periodic spikes with 300 time step intervals, corresponding to a theoretical angular frequency of $\omega^{\prime}=2.09$ $rad/s$, and indeed a cluster at about 2 $rad/s$ is present in the $\omega$-$b_{c}$ plot. The theoretical and optimized angular frequencies aligned, demonstrating learning of the dynamical signal by the network.
@@
Figure 7: Test accuracy vs pruning probability for (a) S-MNIST and (b) PS-MNIST. 1.0 pruning indicate no recurrent weights.
We explored the effect of pruning recurrent weights in an ablation study for the S-MNIST and PS-MNIST tasks between comparable BRF and ALIF networks, as detailed in Table 4 and Table 4. Results are shown in Figure 7, where the BRF neuron provides constant performance throughout pruning, while the ALIF model performs poorly when lacking recurrent weights. This demonstrates that explicit recurrencies were not necessary for the BRF networks with resonating properties to learn well, while these connections were crucial for the ALIF networks. Moreover, the BRF networks’ accuracy had a smaller standard deviation throughout compared to the ALIF networks, also showing that the BRF network performance was not affected by the specific pruned weights as much as the ALIF networks.

### v2

@@ 1 Introduction
Artificial neural networks (ANNs) have become the main method for solving machine learning problems in recent years (Goodfellow et al., 2016; Wu & Feng, 2018; Abiodun et al., 2019). However, they require massive computation and energy, making them inefficient for large-scale real-world applications—particularly in the realm of edge computing— due to the deep non-linear continuous estimators to represent features of the data (Pfeiffer & Pfeil, 2018).
@@
Spiking neural networks (SNNs) circumvent this drawback by processing information through the precise timing of action potentials, or spikes. SNNs can potentially be more efficient compared to ANNs due to their event-driven property, for which computation is only required when spikes are propagated in the system (Paugam-Moisy & Bohté, 2012). Furthermore, SNNs are self-recurrent, as the neurons have a dynamic internal state that modulates their activity over time, and are more biologically realistic than conventional ANNs. Some examples of spiking neurons with ascending biological plausibility include the leaky integrate-and-fire (LIF) neuron, the adaptive leaky integrate-and-fire (ALIF) neuron (Bellec et al., 2018), the Izhikevich neuron (Izhikevich, 2003), and the Hodgkin-Huxley (HH) neuron (Hodgkin & Huxley, 1952). While the HH model is too computationally expensive for practical application, it exhibits rich and complex membrane dynamics that simpler models lack.
@@
One mode of dynamics gaining interest in ML applications in particular is the resonating behavior seen in oscillatory neurons. Subthreshold oscillations of the membrane potential have been observed in various mammalian nervous systems, including the neurons in the frontal cortex (Llinas et al., 1991), in the thalamus (Pedroarena & Llinás, 1997), and in layer II of the medial entorhinal cortex (Alonso & Llinás, 1989), which is crucial in spatial information processing (Doeller et al., 2010). The resonate-and-fire (RF) neuron proposed by Izhikevich (2001) models such damped or sustained subthreshold oscillations of biological neurons: the RF neuron fires when the frequency of the incoming spikes matches that of the neuron’s damped oscillation. Incoming signals with higher frequency leads to more firing for a LIF neuron but less firing for an RF neuron with slow oscillation. As the RF neuron is similarly computationally efficient as the LIF neuron, it is a spiking neuron model potentially suitable for large-scale SNNs (Izhikevich, 2001).
@@
Previous works showed RF neurons implemented in the Intel Loihi 2 (Davies et al., 2018), a neuromorphic processor, can be used to compute the short-time Fourier transform (STFT) of signals (Frady et al., 2022; Shrestha et al., 2023). Moreover, the RF neurons successfully converted raw data signals to spike trains (Shaaban et al., 2024), and further input to SNNs with LIF neurons for detection and classification tasks (Hille et al., 2022; Lehmann et al., 2023). RF neurons have also been implemented in the framework of (feedforward) SNNs as harmonic oscillators for image classification tasks (AlKhamissi et al., 2021), as well as in optical flow estimation and audio classification tasks (Frady et al., 2022). Still, the performance of the RF models did not significantly exceed that of deep LIF SNNs (Frady et al., 2022) or LSTM cells (AlKhamissi et al., 2021), and these studies did not compare with state-of-art spiking neuron models, such as ALIF. Considering the strengths of the RF neuron, these results suggest that the full potential of the RF neuron has yet to be explored.
@@ 2 Resonate-and-Fire Neurons
Refer to captionunstable ($\omega{=}10$, $b{=}-0.3$)stable ($\omega{=}10$, $b{=}-1$)
@@ 3 Balanced RF Models
The influence of implementing the refractory period and the smooth reset can be seen in Figure 2. The single RF neuron was simulated with $\omega=10,\,b=-1$, and the input signal was frequency-timed. The combination of refractory period and smooth reset significantly reduced the number of spikes, while the output spikes effectively reflected the period of the angular frequency of the neuron. The effect of implementing both are highlighted in Figure 10 in Section A.8.
@@
To alleviate the divergence problem, we propose to restrict the parameter space to a subspace below the divergence boundary—an analytically derived relation between $\delta$, $b_{c}$, and $\omega$ that ensures convergence. For an RF neuron to converge or show sustained oscillation, the magnitude of the membrane potential decreases or stays constant with time given an incoming signal:
@@
Furthermore, the explicit form of the subthreshold oscillatory behavior of the membrane potential after an incoming spike is derived in Section A.3 and given as:
@@
With a default $\delta=0.01$, the highest frequency that the neuron can resonate at is the frequency corresponding to an angular frequency of $100\,rad/s$. We define this angular frequency as the upper boundary $\omega_{ub}$. The upper bound of $b_{c}$ that leads to sustained oscillation is implemented in this paper as $p(\omega)$, also considered the divergence boundary:
@@
combined with a trainable b-offset $b^{\prime}>0$ to ensure flexibility and convergence: $b_{c}=p(\omega)-b^{\prime}$, which is constant throughout one sequence length. Figure 9 shows exemplary divergence boundaries for various $\delta$ values. By combining the derived dampening factor with the smooth reset, we present the final equation of $b(t)$ applied for optimization:
@@
Implementing the refractory period, smooth reset, and the divergence boundary leads to efficient learning and sparse spiking for all four datasets.
@@
The frequency response plots shown in Figure 3 assess how well the RF neuron responds to specific frequencies. A peak on a frequency response plot indicates a high sensitivity of the neuron to the respective input signal frequency (details in Section A.4). The figure shows alignment of the response peak and $\omega$ of the RF neuron, demonstrating the resonance property to also be present in the discrete case. The slight offset of the response peak and $\omega$ of the RF neuron is an artifact of the numerical integration of the membrane dynamics’ differential equation and disappears with decreasing the temporal step size $\delta$.
@@
The RF neuron can also be considered a narrow band pass filter, for which only a specific range of frequencies are filtered depending on its angular frequency $\omega$. Combined with the discrete time scale $\delta$ and dampening parameter $b^{\prime}$, they determine how narrow and sensitive the band pass filter is. It should be noted that RF neurons are naturally sensitive to the lower order subharmonics of their angular frequency, that is (with decaying sensitivity) $\frac{1}{2}\omega,\frac{1}{3}\omega,\frac{1}{4}\omega$ and so forth.
@@ 4 Network Implementation
We implement the RF, BRF, and BHRF neurons within RSNNs and applied them to simulations with several benchmark datasets [^1] Source code avaliable at https://github.com/AdaptiveAILab/brf-neurons. The essential formulations of the BRF and the BHRF neuron are summarized in Algorithm 1 and Algorithm 2, respectively. Note that for the algorithmic formulation we change the notation to time-discrete tensor operations and superscript the time index, whereas the transition from $t$ to $t+1$ is considered as a time delay of $\delta$.
@@
We trained the networks with BPTT with the double-Gaussian function (Equation 39, Yin et al., 2021) as the surrogate gradient (Bellec et al., 2018; Neftci et al., 2019). Further details of the networks are described in Section A.6.
Figure 4: Overview of the datasets. Examplary MNIST image and its corresponding sequential and permuted representations. Common pixel row outlined in red on MNIST and S-MNIST sample. ECG sample after level cross encoding. SHD sample after preprocessing.
@@ 5 Experiments
The MNIST dataset consists of grayscale 28$\times$28 pixel hand-written digit images for classification. The sequential-MNIST (S-MNIST), which converts the image to a sequence of 1 $\times$ 784, is a prominent benchmark dataset that enable comparison between sequential models with 54,000 images for training, 6,000 for validation, and 10,000 for testing. In the permuted S-MNIST variant (PS-MNIST), the pixel positions were first permuted randomly and then input to the network sequentially.
Electrocardiogram (ECG) recordings of the heart are represented in voltage over time and consist of cyclic activity with six characteristic waveforms: P, PQ, QR, RS, ST, and TP. The QT database consisted of ECG recordings with per-time step labels classified by experts in the field (Laguna et al., 1997). There were 105 recordings, each recorded for 15 minutes with two electrodes. We left out 24 signals, as they were sudden death recordings without annotation. For the preprocessing shown in Figure 4, refer to Section A.7.
@@ 6 Results
Task Model Architecture No. params. ($\downarrow$) Test Acc. ($\uparrow$) SOPs ($\downarrow$) SOPs/step ($\downarrow$) SOP Ratio ($\downarrow$)
@@
RF 1,256,10 68,874 98.0±$\pm 0.4\,\%$ 29,034.8 37.03 0.41$\times$
BRF 1,256,10 68,874 99.0±$\pm 0.1\,\%$ 15,462.6 19.72 0.22$\times$
BHRF 1,256,10 68,874 99.1±$\pm 0.1\,\%$ 21,565.7 25.51 0.30$\times$
@@
RF 4,36,6 1,734 85.5±$\pm 0.7\,\%$ 11981.9 9.22 0.34$\times$
BRF 4,36,6 1,734 85.8±$\pm 0.7\,\%$ 6,307.7 4.85 0.18$\times$
BHRF 4,36,6 1,734 87.0±$\pm 0.4\,\%$ 6,233.8 4.80 0.18$\times$
SHD DCLS-D** 700,$(256)^{2}$,20 $\approx$200,000 95.1±$\pm 0.2\,\%$
@@
RF 700,128,20 108,820 89.2±$\pm 0.6\,\%$ 4750.2 19.00 0.19$\times$
BRF 700,128,20 108,820 91.7±$\pm 0.8\,\%$ 3,502.6 14.01 0.14$\times$
BHRF 700,128,20 108,820 92.7±$\pm 0.7\,\%$ 4,139.5 16.56 0.17$\times$
The results, the network architecture, the number of parameters, and the number of produced spikes between the best-performing BRF-, RF- and BHRF-RSNN and the baseline ALIF-RSNNs (see Section A.2) or other state-of-the-art (SoTa) models, are compared for each classification task in Table 1. The BRF- and BHRF-RSNN hyperparameters for each dataset are shown in Table 3 in Section A.12. The BHRF and BRF models outperformed the RSNN SoTa for four and three tasks respectively. Complementing accuracy, we also find much enhanced sparseness in the BRF and BHRF networks while using substantially fewer parameters compared to the ALIF networks. Note that RF model shown in Table 1 did not include a reset. Adding either a hard or soft reset to the vanilla RF neuron dramatically affects its task performance as shown in Figure 11 in Section A.9.
We also compare the theoretical energy efficiency of the RF and ALIF models by computing their respective total spiking operations (SOPs) (see Equation 41) and SOPs per sequence step. As both measures were considerably smaller for the BRF and BHRF models, this implies they require less computation to achieve better or comparable performance. This may be due to the resonating spiking behavior of the neuron, needing fewer spikes to represent its frequency, as well as the refractory period and smooth reset that hinders continuous spiking. A particularly large difference was seen in the SHD task, where the BRF model spiked only 14% of the ALIF model.
Compared to the standard RF model without reset, our balanced variants achieved significantly sparser activity while obtaining higher performance. Especially for the PS-MNIST, the standard RF model stayed at chance, while our balanced RF variations outperformed RSNN-SoTA. This is due to the damping factor and angular frequency combination leading to a highly diverging behavior. It especially affects the PS-MNIST dataset, as it requires a wider range of resonant frequencies to counter the increased randomness in the signal. For example, neurons with $\omega=70$ rad/s and $b_{c}=-1$ leads to higher magnitude towards the end of the sequence. Having many of such neurons causes the whole system to diverge and become unstable. On a different note, the performance of the BRF model depends on the initialization of the parameters and it generally under-performs when the angular frequencies are too far from the frequencies underlying the dataset.
Overall, we find that RSNNs comprised of either BRF and BHRF neurons exceeded performance for all classification tasks compared to the baseline ALIF-RSNN. Compared to standard RF neurons, the smooth reset, refractory period, and divergence boundary significantly improved the stability and efficiency of the RF parameters, possibly exploiting the model’s resonant properties. We study this in detail below.
@@
Task Model $\mathbf{95\,\%}$ ($\downarrow$) $\mathbf{98\,\%}$ ($\downarrow$) $\mathbf{100\,\%}$ ($\downarrow$)
S-MNIST ALIF 105 162 276
RF 134 172 263
BRF 3 29 246
BHRF 5 12 119
PS-MNIST ALIF 143 200 265
BRF 10 39 282
BHRF 6 19 123
ECG ALIF 51 157 282
RF 8 29 112
BRF 6 32 75
BHRF 15 38 109
SHD ALIF 14 15 16
RF 2 4 5
BRF 2 3 7
BHRF 3 5 8
@@
Figure 5: Top row: S-MNIST, PS-MNIST, ECG, and SHD learning curve between BRF, RF and ALIF model. Each curve averaged per epoch (solid line) with standard deviation (shaded area) over five runs. The dot on the accuracy curves depict the point at which 95 % of the final accuracy was reached. Bottom row: S-MNIST, PS-MNIST, ECG, and SHD initial and optimized BRF parameter combinations: angular frequency $\omega$ and b-offset $b^{\prime}$ for all runs. Dashed line is the divergence boundary. RF models are simulated without reset. For convergence with traditional reset see Section A.9.
The RF, BRF, and ALIF model learning curves are presented in Figure 5. We reduced the ALIF model to the same size and number of trainable parameters as the BRF model. For the S-MNIST and PS-MNIST tasks, the ALIF networks only learn effectively when the truncated BPTT (TBPTT) with truncation step 50 is applied. Nonetheless, the fully back-propagated BRF networks converge significantly faster than the TBPTT ALIF networks, also evident from the quantitative convergence results (Table 2).
A similar pattern of fast and stable convergence is observed for the ECG dataset and SHD. The convergence between RF and BRF for SHD and ECG are similar because the vanilla RF neurons resonate with lower frequencies, affected less by diverging behavior. Note that the RF model shown in Figure 5 did not include a reset. Adding either soft or hard reset leads to slower and more unstable learning, as shown in Figure 11 in Section A.9. Figure 5 highlights the variation in training of RF networks depending on the dataset, which is alleviated by the BRF network.
While numerically large, the learning rates of the (B)RF-RSNNs led to stable and fast convergence for most datasets. This is in contrast to standard SNNs where large learning rates lead to poor performance. Here, the double-Gaussian surrogate gradient function (Yin et al., 2021) we used may have contributed by ensuring gradient flow even for neurons far from firing. Combined with the high learning rate, this may have enabled non-spiking neurons to effectively shift their angular frequencies.
The ECG task was learned most effectively for the BRF and BHRF networks with a small batch size and high learning rate. A smaller batch size can better capture detailed variations within the dataset, notably the timing of the input spikes, which may have been important for per-step wave classification.
Ongoing investigations (Higuchi et al., 2024) suggest that the shape of the error landscape may be a major contributor to the fast and stable convergence of the BRF model (see Figure 13 in Section A.11).
@@
For the S-MNIST task, plotting $\omega$-$b_{c}$ (Figure 5) shows clusters around $\omega$ values of 18 to 28 $rad/s$ and around $\omega$ of 40 $rad/s$ with a wide range of $b_{c}$. The dampening factor close to zero indicates near-sustained oscillation behavior, preserving resonance amplitude for an extended duration. The clustered variety in $b_{c}$ values suggests that short- and long-term memory are essential for effective learning of the S-MNIST task in B(H)RF networks. Additionally, the distribution of $\omega$ for S-MNIST task showed a bimodal distribution with the highest peak at 23.3 to 25.3 $rad/s$ and the second peak at 40.8 to 42.8 $rad/s$.
The MNIST digits all contain some form of diagonal line with a width of 3 to 6 pixels, most prominently observed in digit 1. When converted into a sequence row-by-row, the lines are distributed among the sequence with a periodic signal of about 25 pixels. Considering the pixels as discrete time steps of $\delta$ = 0.01, the theoretical period T frequently observed in the signal is T = 0.25 $s$. Thus, the most frequent theoretical angular frequency $\omega^{\prime}$ underlying the signal is around $\omega^{\prime}=2\pi/0.25\,s=25.13\,rad/s$ which is close to the most frequently learned angular frequencies of the BRF neuron. This simple calculation thus suggests that the BRF neurons were able to learn meaningful frequencies that underlie the S-MNIST dataset. For the PS-MNIST, we see more spread of $\omega$ and $b_{c}$ parameters, as the specific characteristic frequency structure from S-MNIST is essentially obscured by the permutation.
Inspecting spike train ECG samples (Figure 4), we observe distinguishable periodic spikes with 300 time step intervals, corresponding to a theoretical angular frequency of $\omega^{\prime}=2.09$ $rad/s$, and indeed a cluster at about 2 $rad/s$ is present in the $\omega$-$b_{c}$ plot. The theoretical and optimized angular frequencies align, demonstrating learning of the dynamical signal by the network.
@@
Figure 7: Test accuracy over pruning probability for (a) S-MNIST and (b) PS-MNIST. 1.0 pruning indicate no recurrent weights.
We explored the effect of pruning recurrent weights in an ablation study for the S-MNIST and PS-MNIST tasks between comparable BRF and ALIF networks, as detailed in LABEL:app:smnist_lc_hyp and results are shown in Figure 7. While the BRF neuron provides consistent performance throughout pruning, the ALIF model performs poorly when lacking recurrent weights. This demonstrates that explicit recurrencies are not necessary for the BRF networks with resonating properties to learn reasonably well, while these connections are crucial for the ALIF networks. Moreover, the BRF networks’ accuracy has a smaller standard deviation throughout compared to the ALIF networks, also showing that the BRF network performance is not affected by the specific pruned weights as much as the ALIF networks.